OpenStudy (anonymous):

The fizz produced when a tablet of Alka-Seltzer dissolves in H2O comes from the reaction of sodium carbonate,, NaHCO3, and citric acid, C6H8O7: 3NaHCO3+C6H8O7--> 3CO2+3H2O+Na3C6H5O7 supposed the tablet containing 1.00g of sodium bicarbonate and 1.00g of citric acid reacts to completion. How many grams of CO2 are produced and how many grams of the excess reagent remain when the reaction is complete?

OpenStudy (matt101):

First we have to figure out the limiting reagent, so calculate moles (mass/molar mass). moles of sodium bicarbonate = 1/84 = 0.0119 moles of citric acid = 1/192 = 0.0052 Divide moles by the coefficient in the equation to see which is limiting: Sodium bicarbonate: 0.0119/3 = 0.0040 Citric acid: 0.0052/1 = 0.0052 Therefore sodium bicarbonate is limiting. Since the coefficients of sodium bicarbonate and CO2 are the same in the equation, the number of moles of sodium bicarbonate consumed equals the number of moles of CO2 produced. Therefore 0.0119 moles of CO2 are produced, which is 0.524 g (after multiplying by molar mass). According to the equation, 3 moles of sodium bicarbonate react with 1 mole of citric acid. That means you have a third as many moles of citric acid as sodium bicarbonate, which in this case is 0.119/3 = 0.0040 moles. However, we know from above that there were 0.0052 moles to begin with. Therefore, after the reaction, there are 0.0052-0.0040=0.0012 moles of citric acid remaining. Multiplying by the molar mass of citric acid, we get 0.23 g remaining.

OpenStudy (anonymous):

Thanks Matt!! That's I got as well; except I left the second part as 0.238g being that we have three sig. figs. from the 1.00g